Exercise 5-1

On January 1, 2013, Pam Company purchased an 85% interest in Shaw Company for $541,800. On this date, Shaw Company had common stock of $398,100 and retained earnings of $143,700.

An examination of Shaw Company’s assets and liabilities revealed that their book value was equal to their fair value except for marketable securities and equipment:
Book Value Fair Value
Marketable securities $19,800 $44,600
Equipment (net) 120,200 140,000

(a)

Prepare a Computation and Allocation Schedule for the difference between book value of equity acquired and the value implied by the purchase price. (Round answers to 0 decimal places, e.g. 5,125.)Marketable securities Equipment (net) Book Value Fair Value $19,800 $44,600 120,200 140,000 (a) Prepare a Computation and All
Answer #1

Computation and Allocation Schedule for the difference book value of equity aquired and the value implied by the purchase price

85% 15% 100%
Parent share NCI share Equity value
Purchase Value and implied value(A) $541,800 $95,611 $637,411
Book Value of Equity acquired:
Commom Stock(B) 338,385 59,715 398,100
Retained earnings(C) 122,145 21,555 143,700
Total Book value(D)=B+C 460,530 81,270 541,800
Difference between implied and book value (E)=A-D 81,270 14,341 95,611
Marketable Securities (F) (21,080) (3,720) (24,800)
Equipment (G) (16,830) (2,970) (19,800)
Balance (E-F-G) 43,360 7,651 51,011
Record New Goodwill (43,360) (7,651) (51,011)
Balance $0 $0 $0

Note: (i) Purchare value and implied value=15% NCI value= $541,800\times100/85\times15/100=$95,611

(ii) Marketable Securties = 100% Equity Value=Fair Value - Book Value=$44,600-$19,800=$24800

(iii) Equipment = 100% Equity Value = Fair value- Book value = $140,000-$120,200=$19800